OpenStudy (anonymous):

Find the value of 3x√x^3

OpenStudy (anonymous):

i got 3x√x over x. is that correct?

OpenStudy (anonymous):

no...its not. show me what u did.

OpenStudy (anonymous):

first i simplified the √x^3 and got x^2√x

OpenStudy (anonymous):

then multiplied the top and bottom by √x to get rid of the radical in the denominator. so i have 3x√x over x^2. but my options don't even have that so i went with 3x√x over x

OpenStudy (anonymous):

that's where u went wrong...look here: \[\sqrt{x ^{3}} = \sqrt{x ^{2}x} = x \sqrt{x}\] this is what u should get after simplification.

OpenStudy (anonymous):

yep. forgot you had to split them first. thank you

OpenStudy (anonymous):

sure, no prob!

OpenStudy (anonymous):

still can't find the answer

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

why not? multiply what i got earlier with 3x.

OpenStudy (anonymous):

well I'm dividing 3x by x√3 so its not multiplying

OpenStudy (anonymous):

wait...ur question is \[3x / x \sqrt{3} ?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh! i got it wrong! my bad! well...then first of all, cancel out the x's. then rationalize the denominator.

OpenStudy (anonymous):

so 3/x?

OpenStudy (anonymous):

no...there won't be any x in the answer.

OpenStudy (anonymous):

all of my options have x in them...

OpenStudy (anonymous):